OpenStudy (anonymous):

Base 7= B7 Evaluate the log without a calculator logB7 6=0.9 logB7 4=0.7 logB7 5=0.8 Find logB7 28/5

OpenStudy (anonymous):

@mathmale

OpenStudy (isaiah.feynman):

You want everything to solved for you. While you watch.

OpenStudy (anonymous):

no, i just need to know how to do it

OpenStudy (isaiah.feynman):

Base 7=B7? Is that the original question?

OpenStudy (anonymous):

yes @Isaiah.Feynman

OpenStudy (triciaal):

28 can be expressed as 4*7 ; 5 can be expressed as (6 + 4) / 2 when you multiply numbers add the logs when you divide numbers subtract the logs this should help

OpenStudy (anonymous):

so then i put logB7 (4*7)- logB7 (6+4/2)?

OpenStudy (mathmale):

Hello, Shay! Glad to see you again! logB7 6=0.9 logB7 4=0.7 logB7 5=0.8 Find logB7 28/5 Here's what I'd do: Use the rule of logs for division: log (a/b) = log a - log b. Therefore, logB7 (28/5) = logB7 (28)-logB7 (5) Note that logB7 (5) is given (above), but that logB7 (28) is not. However, logB7 (28) can be written as logB7 (4) + logB7 (7). logB7 (7) is not given above. However, if you recognize that logB7 (7^1) = 1, the rest is easy. Note how ln e^1 = ln e = 1; log 10^5=5, and so on. Ask me if this is confusing for you. so, now let's put everything together. logB7 (28/5) = logB7 (4) + logB7 (7) - logB7 (5). Please review what I've typed above and then try to arrive at a final answer. Show your work, so that I can give you suitable feedback.